OpenStudy (anonymous):

The table below shows four systems of equations: Which pair of systems will have the same solution?

OpenStudy (triciaal):

cannot see the table.

OpenStudy (anonymous):

System 1 System 2 System 3 System 4 4x − 5y = 2 4x-5y=2 4x - 5y =2 4x-5y=2 3x − y = 8 3x − 8y = 4 13x − 8y = 26 10x + 3y = 15

OpenStudy (anonymous):

System 1 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 1 to two times the second equation in system 1 System 2 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 2 to two times the second equation in system 2. System 2 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 2 to three times the second equation in system 2. System 1 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 1 to three times the second equation in system 1.

jigglypuff314 (jigglypuff314):

mmm I suggest solving systems 1, 2, and 3 first :)

OpenStudy (anonymous):

how do I solve them

OpenStudy (anonymous):

example please

jigglypuff314 (jigglypuff314):

like for system 4 (which you probably won't need since no options mentions it) to solve it by elimination method choose a variable and turn it to LCM like 4x - 5y = 2 --> 12x - 15y = 6 10x + 3y = 15--> 50x + 15y = 75 then add ----------------- so 62x + 0 = 81 62x = 81 then divide both sides by 62 so that x = 81/62 you would then plug it in to one of the original equations :) 10(81/62) + 3y = 15 810/62 + 3y = 15 3y = 15 - 810/62 y = 20/31 so the solution would be x = 81/62 and y = 20/31

OpenStudy (anonymous):

so the answer will be A

OpenStudy (anonymous):

its out of A and D

jigglypuff314 (jigglypuff314):

3x - y = 8 --> multiply through by "3" --> 9x - 3y = 24 plus 4x - 5y = 2 does equal 13x − 8y = 26

jigglypuff314 (jigglypuff314):

therefore the second equation in system 3 is obtained by adding the first equation in system 1 to three times the second equation in system 1.

OpenStudy (anonymous):

thank you can you help me with another its not so complicated like this one

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

before I post it can you translate this word problem to number form x the square root of the quantity x plus 7

jigglypuff314 (jigglypuff314):

x√(x+7)

OpenStudy (anonymous):

now is the x just before the square root or on top of the small line on the line

OpenStudy (anonymous):

on the left

jigglypuff314 (jigglypuff314):

to the left because it did not say "x root" it said "square root"

jigglypuff314 (jigglypuff314):

you still there? I only have a few minutes left before I have to go :)

OpenStudy (anonymous):

Which graph best represents the function y = x the square root of the quantity x plus 7

OpenStudy (anonymous):

I tried to input the equation you put but the graph didn't look like these

jigglypuff314 (jigglypuff314):

then I'm guessing it was y = root (x+7) parent graph was y = root (x) http://www.ltcconline.net/greenl/courses/103a/keys/exam2PracticeExam/img21.jpg how it moves is determined by y = root (x - h) + k (h, k) being the vertex so for y = root (x - (-7)) ^ you'll get (-7, 0) as the vertex

OpenStudy (anonymous):

so its the second graph

OpenStudy (anonymous):

wait nvrm the last one

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

I don't want to be a burden but do you have time with another ill start new post so I csn give medal

jigglypuff314 (jigglypuff314):

It wouldn't be a burden, but I've really got to go, sorry :/ perhaps you can tag esshotwired or inkyvoyd or SolomonZelman or shamil98 to help you with your question :) just put at @ in front of their name to tag them like @masterchief609 good luck!

OpenStudy (anonymous):

ok thanks for all the help really appreciate it

jigglypuff314 (jigglypuff314):

you're welcome :)